A proton is accelerated from rest through a potential difference of 5000 v. What is the minimum uncertainty in the position of the proton if the uncertainty in its momentum is 3%? (Use ∆x∆p = ℏ)
Answer #1

q\Delta V=\frac{mv^{2}}{2}=\frac{(mv)^{2}}{2m}=\frac{p^{2}}{2m}

p=\sqrt{2mq\Delta V}

\frac{\Delta p}{p}*100=3

\Delta p=\frac{3*\sqrt{2mq\Delta V}}{100}

\Delta x=\frac{h*100}{2\pi *3*\sqrt{2mq\Delta V}}=\frac{6.626*10^{-34}*100}{2\pi *3*\sqrt{2*1.67*10^{-27}*1.6*10^{-19}*5000}}

\Delta x=2.15*10^{-12}\ m